# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v1

@@ Abstract
Quantum entanglement between the spin, orbital and lattice degrees of freedom in condensed matter systems can emerge due to an interplay between spin-orbit and vibronic interactions. Heavy transition metal ions decorated on a face-centered cubic lattice, for example in $5d^{1}$ double perovskites, are particularly suited to support these quantum entangled states, but direct evidence has not yet been presented. In this work, we report additional peaks in the low-energy spectra of a $5d^{1}$ double perovskite, Ba2CaReO6, which cannot be explained by adopting a purely classical description of lattice vibrations. Instead, our theoretical analysis demonstrates that these spectroscopic signatures are characteristic of orbital-lattice entangled states in Ba2CaReO6. Crucially, both theory and experiment demonstrate that these quantum-entangled states persist to low temperatures, despite the onset of multipolar order.
@@
This symmetry-allowed coupling finds experimental support in $5d^{1}$ DP compounds such as Ba2NaOsO6 [4, 5, 6], Cs2TaCl6 [7, 8], and Ba2MgReO6 [9, 10] where structural orderings have been observed. A recent study using synchrotron elastic x-ray scattering techniques further confirm these findings by unveiling the simultaneous development of quadrupolar and structural orders [10]. Complementing these experimental observations are first-principles calculations, which have successfully replicated quadrupolar orders coupled with cooperative JT deformations [11, 12].
However, these previous studies of $5d^{1}$ DPs have predominately favored a classical description of the lattice degrees of freedom, neglecting the quantum nature of lattice vibrations which play a decisive role in shaping the low-energy spectra. In particular, the strong spin-orbit coupling in $5d$ ions significantly reduce the potential energy barrier between different types of static octahedral deformations. As such, the lattice of $5d^{1}$ DPs can exist in a dynamic superposition of distinct static JT distortions, reflecting its inherent quantum nature [13].
@@
Crucially, these $5d^{1}$ DPs should exhibit the presence of vibronic states (derived from portmanteau of vibrational and electronic), resulting from the quantum entanglement between electronic $d^{1}$ states on metal sites with the quantized vibrational modes of the surrounding lattice. The experimental fingerprints of these orbital-lattice entangled states are additional transitions, which emerge due to the dressing of the ground ($j_{\mathrm{eff}}$=3/2) and excited ($j_{\mathrm{eff}}$=1/2) manifold by lattice vibrations that cannot simply be explained by adopting a purely classical description.
@@
Indeed, recent investigations hint that these quantum effects play a significant role in $5d^{1}$ DPs. For example, recent resonant inelastic x-ray scattering (RIXS) measurements suggest the presence of the dynamic JT effect in several Re-based compounds [14], where the fine structures observed in the spectra were theoretically reproduced by assuming the development of the orbital-lattice entanglement in an undistorted ReO6 octahedron. Further supporting this notion are first-principles calculations of several $5d^{1}$ DPs, which indicate that the dynamic JT effect plays a non-negligible role in stabilizing the ground state [15]. Moreover, the emergence of multipolar phases observed in the wider family of cubic $5d^{1}$ DP compounds can trace their natural origin to the ordering of vibronic states at sufficiently low temperatures [16]. Nevertheless, several open questions remain unanswered.
@@
For instance, there is a notable absence of direct spectroscopic evidence for vibronic states in $5d^{1}$ DPs. This is because the RIXS experiments conducted so far have primarily been focused on the metal absorption $L_{3}$-edge, resulting in resolution-limited spectra that fail to capture the detailed low-energy features of vibronic states. To conclusively demonstrate the presence of vibronic dynamics, RIXS experiments with higher energy resolution are imperative. One promising avenue is to investigate RIXS at the absorption $K$-edge of the ligands, which are strongly covalently bonded to the metal ions. RIXS at the O $K$-edge is particularly sensitive to vibronic excitations due to its significantly longer core-hole life time ($\Gamma$=150 meV) [17].
@@
To understand how vibronic states arise in the first place, we review the key onsite interactions of a $d^{1}$ ion residing in an undistorted octahedral environment. Here, the covalency between the metal $d$ and ligand $p$ orbitals splits the five atomic $d$ orbitals into doubly-degenerate $e_{g}$ and triply-degenerate $t_{2g}$ orbitals. Given that the $t_{2g}$ orbitals are lower in energy, they are populated by the single $d$ electron.
@@
via $g$, the dimensionless vibronic coupling parameter between the $d$-$p$ hybridized electronic orbitals $|\gamma\rangle$ ($\gamma=yz,zx,xy$) and the lattice distortion modes $\hat{q}_{\gamma}$ ($\gamma$=$x^{2}$, $y^{2}$, $z^{2}$) with vibration frequency of $\omega$. Here, $\hat{p}_{\gamma}$ is the dimensionless conjugate momenta. Throughout the work, we consider only the $E_{g}$ vibrational modes [Fig. 1(a)] among the vibronically active modes because their coupling strength is the most dominant.
The resultant adiabatic potential energy surface (APES) of Eq. (3) – where the effects of spin-orbit coupling are neglected – comprises three disconnected harmonic oscillator potentials [Fig. 1(b)]. In such cases, due to the large energy barriers between the degenerate ground states, a purely classical description of the lattice distortion would suffice. However, in cases involving heavy ions (e.g. $5d$ ions) where the effect of spin-orbit interaction becomes important, the potential barrier can be reduced significantly which qualitatively changes the quantum states of the system.
@@
$\hat{V}_{\text{JT}}$ $=\frac{\hslash\omega g}{2}\begin{pmatrix}0&\sqrt{2}\hat{q}_{x^{2}-y^{2}}&-\sqrt{2}\hat{q}_{z^{2}}\\ \sqrt{2}\hat{q}_{x^{2}-y^{2}}&-\hat{q}_{z^{2}}&\hat{q}_{x^{2}-y^{2}}\\ -\sqrt{2}\hat{q}_{z^{2}}&\hat{q}_{x^{2}-y^{2}}&\hat{q}_{z^{2}}\\ \end{pmatrix},$ (4)
@@
Qualitatively, this results in an APES with a shallow barrier between adjacent minima, as shown in Fig. 1(c). Consequently, the wave function spreads over the circular trough of the APES around the conical intersection. The energy eigenstates (vibronic states) of the Hamiltonian $\hat{H}=\hat{H}_{\text{SO}}+\hat{H}_{\text{JT}}$ take the orbital-lattice entangled form,
@@
Figure 2: The calculated RIXS spectrum at the (a) rhenium $L_{3}$ and (b) oxygen $K$ edge based on a classical ($g$=0) and quantum ($g$=1.8) description of the lattice degrees of freedom are denoted by the dashed blue and solid red lines, respectively. The faint lines correspond to calculations for intermediate coupling strengths $g$=0.3, 0.6, 0.9, 1.2 and 1.5.
To illustrate the spectroscopic signatures of these vibronic transitions, we plot in Fig. 2 the calculated RIXS spectra for a range of coupling strengths ($g$), for an undistorted Ba2CaReO6. We employed the Kramers-Heisenberg formula [22, 23] and the dipole approximation to calculate the spectra that arise from transitions between the vibronic states (See for computational method Ref. [14]). We applied the fast collision approximation [24, 25] for Re $L_{3}$-edge spectra due to the short lifetime in the intermediate states [26]. We explicitly considered the intermediate $d^{2}$ states with Hund coupling of 0.26 eV [27] for O $K$-edge spectra.
In the classical regime ($g$=0), our calculations reveal a single, purely-electronic $j_{\mathrm{eff}}$=1/2 transition at $\Delta E$$\sim$0.5 eV, which is characterized by a symmetric peak at both the rhenium $L_{3}$ and oxygen $K$ edges. On the other hand, when a quantum description of the lattice is adopted ($g$=1.8), the $j_{\mathrm{eff}}$=1/2 transition broadens asymmetrically and acquires a long tail that extends towards high energies at both the Re $L_{3}$ and O $K$ edges. This broadening is due to the presence of many additional quantum vibronic transitions, which come from the dressing of the $j_{\mathrm{eff}}$=3/2 manifold by lattice vibrations.
Furthermore, a strong asymmetric peak also appears at $\Delta E$$\sim$ 0.15 eV of the O $K$ RIXS spectra for $g$=1.8 [Fig. 2(b)]. The peak originates from a multitude of quantum vibronic transitions between 0-0.4 eV, which appear due to the dressing of the $j_{\mathrm{eff}}$=3/2 ground states by the lattice degrees of freedom. As shown in Fig. 2(b), such a peak is absent in the classical case, where $g$=0. Therefore, the presence of such a peak is the key spectroscopic fingerprints of quantum vibronic states arising from spin-orbit-lattice entanglement.
Figure 3: Dynamic JT effect in the elastic field. (a)-(c) The APES’s with varied effective elastic field strengths. (d) The evolution of the vibronic energy levels (red dots) with respect to the elastic field arising for tetragonal deformation. $F>0$ ($<0$) corresponds to the tetragonal compression (elongation) in eV. The solid blue and dashed green lines correspond field dependence of the harmonic oscillator frequencies, which are obtained from the parabolic fits depicted on the APES.
@@
The external field alters the shape of the APES and hence modulates the nature of the low-energy vibronic states on sites. For example, under the external field of tetragonal compression type ($F_{z^{2}}>0$), one of the minima of the APES is lowered in energy and shifted towards larger deformations [Fig. 3(a)–(c)]. As the magnitude of $F$ increases, the ground vibronic states split into two Kramers doublets [Fig. 3(d)]. The distribution of the energy spectrum approaches that of a two-dimensional harmonic oscillator around the global minimum of the APES. The solid blue lines in Fig. 3(d) correspond to the two frequencies. Similar behaviour is observed for tetragonal elongation too ($F_{z^{2}}<0$), where the two frequencies are denoted by the dashed green lines.
Next we consider whether the spin-orbit-lattice entanglement is robust against the elastic field in the case of $5d^{1}$ DP materials. By adopting realistic interaction parameters for the rhenium-based DPs of $\lambda=0.32$ eV, $g=1.3$ [14], and $\omega=66$ meV [28], we find that the ground vibronic state is stabilized by 30 meV, which is more than two times larger than the static JT stabilization energy of 14 meV and is three times larger than the effective field of the Re compounds measured by the structural transition temperatures. Therefore, given that the size of $F$ is very small ($\sim$ 10 meV), our calculations indicate that the vibronic entanglement effect in cubic $5d^{1}$ DPs should persist strongly, even below the structural transition temperature.
@@
We calculated the Re $L_{3}$-edge RIXS spectra by adopting the interaction parameters described above and also accounting for the experimental configuration shown in the insert of Fig. 4(a). Our calculations reveal that the vibronic coupling modulates the peak shape of the excited $j$=$1/2$ multiplet state at 0.52 eV considerably, resulting in additional transitions at 0.59, 0.67, 0.73 eV. The calculated transition energies and intensities between the vibronic states, denoted by the grey bars in Fig. 4(a), are in good agreement with the experimental data. This nice correspondence between the two is remarkable given that the calculations were based on an undistorted ReO6 octahedron (i.e. for Ba2CaReO6 above $T_{\mathrm{s}}$), in contrast to the measurements which were performed at $T$ = 7 K. Hence, for a more representative calculation of the transitions, we have to account for the effect of the structural distortions on the vibronic level, as depicted in Fig. 3.
Curiously, the size of the structural distortion across many $5d^{1}$ DPs is very subtle, and in some cases remain below the detection limits of x-ray diffraction. This is also true for single crystalline Ba2CaReO6 where the size of the structural distortion is in the order of 0.2%. Nonetheless, even if we adopt the size of distortion found in polycrystalline Ba2CaReO6 [18, 19] (which are approximately two times larger than those found for the single crystalline samples), we find that the changes to the peak shape is very subtle, resulting only in minor changes to the transition energies and intensities, as denoted by the black lines in Fig. 4(a).
@@
Next, when the cooperative structural deformation is accounted for, we also find that the distribution of the intensities of the O $K$-edge RIXS spectra does not change considerably. The black lines in Fig. 4(b), denote the calculated transitions at $T$=17 K, based on the distorted structure described earlier. Our experimental data and calculations both demonstrate clearly that the vibronic states are impervious to the quadrupolar ordering in Ba2CaReO6. Therefore, the orbital-lattice entanglement and hence the dynamic JT effect persist.
Finally, the weak temperature dependence of the RIXS spectra also supports the presence of the dynamic JT effect. The peak shape in the present O $K$-edge RIXS spectra does not change across the quadrupolar transition  [^1] The peak shape of the Re $L_{3}$-edge RIXS spectra of Ba2MgReO6 also does not change across the quadrupolar transition [14]., suggesting a mechanism independent from the temperature. Although one may assume the static JT effect below the quadrupolar transition and assign the low-energy peaks to the transitions between the JT split adiabatic potential surfaces, this scenario cannot explain why the peaks remain above the quadrupolar transition temperature. The dynamic JT mechanism naturally explains all the features of the RIXS spectra, as demonstrated above.
@@
Note added. After the completion of this work, we became aware of a preprint of  [31] which studied the RIXS of Ba2NaOsO6. Their experimental results at the O $K$ is consistent with the RIXS spectrum presented here but their theoretical model fails to fully explain why the $j_{\mathrm{eff}}$=1/2 and 3/2 peaks are asymmetric. In our work here, we demonstrate clearly that the asymmetry arise from vibronic transitions.

### v2

@@ Abstract
Quantum entanglement between the spin, orbital, and lattice degrees of freedom in condensed matter systems can emerge due to an interplay between spin-orbit and vibronic interactions. Heavy transition metal ions decorated on a face-centered cubic lattice, for example, in $5d^{1}$ double perovskites, are particularly suited to support these quantum entangled states, but direct evidence has not yet been presented. In this work, we report additional peaks in the low-energy spectra of a $5d^{1}$ double perovskite, Ba2CaReO6, which cannot be explained by adopting a purely classical description of lattice vibrations. Instead, our theoretical analysis demonstrates that these spectroscopic signatures are characteristic of orbital-lattice entangled states in Ba2CaReO6. Crucially, both theory and experiment demonstrate that these quantum-entangled states persist to low temperatures, despite the onset of multipolar order.
@@
This symmetry-allowed coupling finds experimental support in $5d^{1}$ DP compounds such as Ba2NaOsO6 [4, 5, 6], Cs2TaCl6 [7, 8], and Ba2MgReO6 [9, 10] where structural orderings have been observed. A recent study using synchrotron elastic x-ray scattering techniques further confirms these findings by unveiling the simultaneous development of quadrupolar and structural orders [10]. Complementing these experimental observations are first-principles calculations, which have successfully replicated quadrupolar orders coupled with cooperative JT deformations [11, 12].
However, these previous studies have predominately favored a classical description of the lattice degrees of freedom, which break down for the $5d^{1}$ DPs. For instance, a resonant inelastic x-ray scattering (RIXS) study of $5d^{1}$ DPs at the $L_{3}$ edge finds an asymmetric $j_{\text{eff}}=1/2$ peak [13], which is characteristic of vibrational transitions between different electronic states described by the Franck-Condon principle. Crucially, the asymmetric shape of the $j_{\text{eff}}=1/2$ peak remains unchanged by structural transitions, thereby ruling out the conventional classical mechanisms as an explanation.
@@
The missing piece in the former studies is the quantum nature of lattice vibrations which play a decisive role in shaping the low-energy spectra. In particular, the strong spin-orbit coupling in $5d$ ions enhances the quantum entanglement between the $j_{\text{eff}}=3/2$ states on metal sites, with the quantized vibrations of the surrounding lattice (dynamic JT effect [14]). The experimental fingerprints of these orbital-lattice entangled vibronic states are additional transitions, which emerge due to the dressing of the ground ($j_{\mathrm{eff}}$=3/2) and excited ($j_{\mathrm{eff}}$=1/2) manifold by lattice vibrations that cannot simply be explained by adopting a purely classical description.
@@
Further supporting this notion are first-principles calculations of several $5d^{1}$ DPs, which indicate that the dynamic JT effect plays a non-negligible role in stabilizing the ground state [15, 16]. Moreover, the emergence of multipolar phases observed in the wider family of cubic $5d^{1}$ DP compounds can trace their natural origin to the ordering of vibronic states at sufficiently low temperatures [17]. Nevertheless, several open questions remain unanswered.
@@
For instance, there is a notable absence of direct spectroscopic evidence for vibronic states in $5d^{1}$ DPs. This is because the RIXS experiments conducted so far have primarily been focused on the metal absorption $L_{3}$-edge, resulting in resolution-limited spectra that fail to capture the detailed low-energy features of vibronic states. To conclusively demonstrate the presence of vibronic dynamics, RIXS experiments with higher energy resolution are imperative. One promising avenue is to investigate RIXS at the absorption $K$-edge of the ligands, which are strongly covalently bonded to the metal ions. RIXS at the O $K$-edge is particularly sensitive to vibronic excitations due to its significantly longer core-hole lifetime ($\Gamma$=150 meV) [18]. Indeed, recent O $K$-edge RIXS measurements of $5d^{1}$ DPs reveal low-lying excitations [19, 20], but their analyses do not account for the dynamic JT effect.
@@
To understand how vibronic states arise in the first place, we review the key onsite interactions of a $d^{1}$ ion residing in an undistorted octahedral environment. Here, the covalency between the metal $d$ and ligand $p$ orbitals splits the five atomic $d$ orbitals into doubly-degenerate $e_{g}$ and triply-degenerate $t_{2g}$ orbitals [23]. Given that the $t_{2g}$ orbitals are lower in energy, they are populated by the single $d$ electron.
@@
via $g$, the dimensionless vibronic coupling parameter between the $d$-$p$ hybridized electronic orbitals $|\gamma\rangle$ ($\gamma=yz,zx,xy$) and the lattice distortion modes $\hat{q}_{\gamma}$ ($\gamma$=$x^{2}$, $y^{2}$, $z^{2}$) with vibration frequency of $\omega$. Here, $\hat{p}_{\gamma}$ is the dimensionless conjugate momenta. Throughout the work, we consider only the $E_{g}$ vibrational modes [Fig. 1(a)] among the vibronically active modes because their coupling strength is the most dominant in $5d^{1}$ DPs [15, 13].
The resultant adiabatic potential energy surface (APES) of Eq. (3) – where the effects of spin-orbit coupling are neglected – comprises three disconnected harmonic oscillator potentials [Fig. 1(b)]. In such cases, due to the large energy barriers between the degenerate ground states, a purely classical description of the lattice distortion would suffice. However, in cases involving heavy ions (e.g., $5d$ ions) where the effect of spin-orbit interaction becomes important, the potential barrier can be reduced significantly [24, 25], which qualitatively changes the quantum states of the system.
@@
$\hat{V}^{\prime}_{\text{JT}}$ $=\frac{\hslash\omega g}{2}\begin{pmatrix}0&\sqrt{2}\hat{q}_{x^{2}-y^{2}}&-\sqrt{2}\hat{q}_{z^{2}}\\ \sqrt{2}\hat{q}_{x^{2}-y^{2}}&-\hat{q}_{z^{2}}&\hat{q}_{x^{2}-y^{2}}\\ -\sqrt{2}\hat{q}_{z^{2}}&\hat{q}_{x^{2}-y^{2}}&\hat{q}_{z^{2}}\\ \end{pmatrix},$ (4)
@@
Qualitatively, this results in an APES with a shallow barrier between adjacent minima, as shown in Fig. 1(c) [24, 25]. Consequently, the wave function spreads over the circular trough of the APES around the conical intersection. The energy eigenstates (vibronic states) of the Hamiltonian $\hat{H}=\hat{H}_{\text{SO}}+\hat{H}_{\text{JT}}$ take the orbital-lattice entangled form,
@@
Figure 2: The calculated RIXS spectrum at the (a) rhenium $L_{3}$ and (b) oxygen $K$ edge based on a classical ($g$=0) and quantum ($g$=1.2) description of the lattice degrees of freedom are denoted by the dashed blue and solid red lines, respectively. The faint lines correspond to calculations for intermediate coupling strengths $g$=0.3, 0.6 and 0.9.
To illustrate the spectroscopic signatures of these vibronic transitions, we plot in Fig. 2 the calculated RIXS spectra for a range of coupling strengths ($g$), for an undistorted Ba2CaReO6. We employed the Kramers-Heisenberg formula [27, 28] and the dipole approximation to calculate the spectra that arise from transitions between the vibronic states (see computational details). We applied the fast collision approximation [29, 30] for Re $L_{3}$-edge spectra due to the short lifetime in the intermediate states [31]. We explicitly considered the intermediate $d^{2}$ states with Hund coupling of 0.26 eV [32] for O $K$-edge spectra.
In the classical regime ($g$=0), where the orbital and lattice degrees of freedom are decoupled, our calculations reveal a single, purely-electronic $j_{\mathrm{eff}}$=1/2 transition at $\Delta E$$\sim$0.5 eV. This transition is characterized by a symmetric peak at both the rhenium $L_{3}$ and oxygen $K$ edges. In contrast, when a quantum description of the lattice is adopted ($g>0$), the orbital and lattice degrees of freedom become entangled. For instance, with $g$=1.2, the $j_{\mathrm{eff}}$=1/2 transition broadens asymmetrically and acquires a long tail that extends towards high energies at both the Re $L_{3}$ and O $K$ edges. This broadening is due to the presence of many additional quantum vibronic transitions, which come from the dressing of the $j_{\mathrm{eff}}$=3/2 manifold by lattice vibrations.
Furthermore, a strong asymmetric peak also appears at $\Delta E$$\sim$ 0.15 eV of the O $K$ RIXS spectra for $g$=1.2 [Fig. 2(b)]. The peak originates from a multitude of quantum vibronic transitions between 0-0.4 eV, which appear due to the dressing of the $j_{\mathrm{eff}}$=3/2 ground states by the lattice degrees of freedom and the difference in the dipole transition paths compared with the $L_{3}$-edge RIXS. As shown in Fig. 2(b), such a peak is absent in the classical case, where $g$=0. Therefore, the presence of such a peak is the key spectroscopic fingerprint of quantum vibronic states arising from spin-orbit-lattice entanglement.
Figure 3: Dynamic JT effect in the elastic field. (a)-(c) The APES’s with varied effective elastic field strengths. (d) The evolution of the vibronic energy levels (red dots) with respect to the elastic field arising for tetragonal deformation. The points at $F=0$ correspond to the vibronic excitation energies shown in Fig. 2(b). $F>0$ ($<0$) corresponds to the tetragonal compression (elongation) in eV. The solid blue and dashed green lines correspond field dependence of the harmonic oscillator frequencies, which are obtained from the parabolic fits depicted on the APES.
@@
The external field alters the shape of the APES and hence modulates the nature of the low-energy vibronic states on sites. For example, under the external field of tetragonal compression type ($F_{z^{2}}>0$), one of the minima of the APES is lowered in energy and shifted towards larger deformations [Fig. 3(a)–(c)]. As the magnitude of $F$ increases, the ground vibronic states (denoted by the red dots) split into two Kramers doublets [Fig. 3(d)]. The distribution of the energy spectrum approaches that of a two-dimensional harmonic oscillator around the global minimum of the APES. The solid blue lines in Fig. 3(d) correspond to the two frequencies. Similar behaviour is observed for tetragonal elongation too ($F_{z^{2}}<0$), where the two frequencies are denoted by the dashed green lines.
Next, we consider whether the spin-orbit-lattice entanglement is robust against the elastic field in the case of $5d^{1}$ DP materials. By using realistic interaction parameters for the rhenium-based DPs of $\lambda=0.32$ eV, $g=1.3$, and $\omega=66$ meV [33] (see computational details), we find that the ground vibronic state is stabilized by 30 meV, which is more than two times larger than the static JT stabilization energy of 14 meV and is three times larger than the effective field of the Re compounds measured by the structural transition temperatures. Therefore, given that the size of $F$ is very small ($\sim$ 10 meV), our calculations indicate that the vibronic entanglement effect in cubic $5d^{1}$ DPs should persist strongly, even below the structural transition temperature.
@@
We calculated the Re $L_{3}$-edge RIXS spectra by adopting the interaction parameters described above and also accounting for the experimental configuration shown in the insert of Fig. 4(a). Our calculations reveal that the vibronic coupling modulates the peak shape of the excited $j_{\text{eff}}$=$1/2$ multiplet state at 0.52 eV considerably, resulting in additional transitions at 0.59, 0.67, 0.73 eV. The calculated transition energies and intensities between the vibronic states, denoted by the grey bars in Fig. 4(a), are in good agreement with the experimental data. This nice correspondence between the two is remarkable given that the calculations were based on an undistorted ReO6 octahedron (i.e. for Ba2CaReO6 above $T_{\mathrm{s}}$), in contrast to the measurements which were performed at $T$ = 7 K. Hence, for a more representative calculation of the transitions, we have to account for the effect of the structural distortions on the vibronic level, as depicted in Fig. 3.
Curiously, the size of the structural distortion across many $5d^{1}$ DPs is very subtle, and in some cases remains below the detection limits of x-ray diffraction. This is also true for single crystalline Ba2CaReO6 where the size of the structural distortion is in the order of 0.2%. Nonetheless, even if we adopt the size of distortion found in polycrystalline Ba2CaReO6 [21, 22] (which are approximately two times larger than those found for the single crystalline samples), we find that the changes to the peak shape are very subtle, resulting only in minor changes to the transition energies and intensities, as denoted by the black lines in Fig. 4(a).
@@
Next, when the cooperative structural deformation is accounted for, we also find that the distribution of the intensities of the O $K$-edge RIXS spectra does not change considerably. The black lines in Fig. 4(b), denote the calculated transitions at $T$=17 K, based on the distorted structure described earlier. Our experimental data and calculations both demonstrate clearly that the vibronic states are impervious to the quadrupolar ordering in Ba2CaReO6. Therefore, the orbital-lattice entanglement and, hence, the dynamic JT effect persist.